OpenStudy (anonymous):

Find the perimeter of the rectangle in simplest form. A rectangle with a width of 3 n squared + 7 n and a length of 5 n squared + 6 n + 3. A. 16n2 + 26n + 6 B. 15n2 + 42n + 3 C. 8n2 + 13n + 3

OpenStudy (johnweldon1993):

We know the perimeter of a rectangle is \[\large P = 2L + 2W\] So we know L = 5n^2 + 6n + 3 And we know W = 3n^2 + 7n so we have \[\large P = 2(5n^2 + 6n + 3) + 2(3n^2 + 7n)\] \[\large P = 10n^2 + 12n + 6 + 6n^2 + 14n\] Now just combine like terms...

OpenStudy (anonymous):

Thank you very much!!! Your a life saver!! =)

OpenStudy (johnweldon1993):

No problem :)